# An extension of Viennot’s shadow to rook placements via orbit harmonics

### v2

@@ 4.4. Standard monomial basis and Hilbert series
Proof of Theorem 4.12.
Claim: $I_{n,m,r}\subseteq{\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$.
In fact, for any rook placement ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$, each column (resp. row) of $n]\times[m]$ has at most one rook. Consequently, for all $1\leq i\leq n$ and $1\leq j,j^{\prime}\leq m$ we have $x_{i,j}\cdot x_{i,j^{\prime}}\in{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$ and thus $x_{i,j}\cdot x_{i,j^{\prime}}\in{\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$ (resp. for all $1\leq i,i^{\prime}\leq n$ and $1\leq j\leq m$ we have $x_{i,j}\cdot x_{i^{\prime},j}\in{\mathrm{gr}}{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$). Additionally, since $\mathcal{R}$} has at least $r$ rooks, any union of nr+1n-r+1 distinct columns of $n]\times[m]$ must meet $\mathcal{R}$}. Thus, for all $1\leq i_{1}<\cdots<i_{n-r+1}\leq n$ we have
@@
Similarly, for all $1\leq j_{1}<\cdots<j_{m-r+1}\leq m$ we have
@@
Then Definition 3.1 indicates the claim above.
The claim above yields a surjective $\mathbb{C}$}-linear map
@@
implying that $\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r})\geq\dim_{\mathbb{C}}(R({\mathcal{UZ}}_{n,m,r}))$. However, Lemma 4.10 indicates that $\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r})\leq|{\mathcal{UZ}}_{n,m,r}|=\dim_{\mathbb{C}}(R({\mathcal{UZ}}_{n,m,r}))$. Therefore, we conclude that
@@
where $c_{{\mathcal{R}}^{\prime}}\in{\mathbb{C}}$. However, Remark 4.11 indicates that the right-hand side of Equation 4.2 can be replaced with a linear combination of some extended shadow monomials
@@
Theorem 4.12 immediately gives the Hilbert series of $R({\mathcal{UZ}}_{n,m,r})={\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. Given a permutation $w\in{\mathfrak{S}}_{n+m-r}$, a increasing subsequence of length $k$ is a sequence $1\leq i_{1}<\cdots<i_{k}\leq n+m-r$ of integers such that $w(i_{1})<\cdots<w(i_{k})$. We write
@@
Let $k$ be the number of shadow lines of $w$ that we use to construct ${\mathcal{S}}(w)$. Write $w\mapsto(P,Q)$ according to the Schensted correspondence, where $P,Q$ have the same shape $\lambda$. The shadow line construction in Subsection 4.1 indicates that $k=\lambda_{1}$. In addition, Schensted correspondence possesses the standard property [15, Thm. 1] $\lambda_{1}=\mathrm{lis}(w)$, so we have $k=\mathrm{lis}(w)$ and thus Equation (4.4) holds. ∎
@@ 5. Module structure
For $d,a,b,p,q\in{\mathbb{Z}}_{\geq 0}$, we consider
$F=\sum_{\mu\vdash d}\{s_{\mu}\cdot h_{a}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{b}\}_{\lambda_{1}=q}\in\Lambda\otimes_{{\mathbb{C}}(q)}\Lambda$
@@
Let $d,a,b,p,q\in{\mathbb{Z}}_{\geq 0}$ be nonnegative integers. Then we have
@@
If one of them is not satisfied, Lemma 5.1 indicates that $\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F=0=\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle G$. We henceforth suppose that all these three conditions hold. Then Lemma 5.1 indicates that
@@
Pieri’s rule and Lemma 5.2 indicate that the left-hand side equals
@@
  • $p\geq n-d$ and $q\geq m-d$ together indicate that $n+m-d-\max\{p,q\}\leq\min\{m,n\}$.
@@
Therefore, $0\leq\min\{m,n\}-r^{\prime}=n+m-\max\{m,n\}-r^{\prime}\leq n+m-\max\{p,q\}-r^{\prime}\leq n+m-(\max\{m,n\}-r^{\prime})-r^{\prime}=\min\{m,n\}$, so we have $0\leq n+m-\max\{p,q\}-r^{\prime}\leq\min\{m,n\}$ and hence the first equality under the summation in (5.4) can be discard. Now we suppose that $p\leq q$ without loss of generality. Then (5.6) indicates that $\max\{p,q\}+r^{\prime}-m=q+r^{\prime}-m\leq\mu_{1}$, and (5.5) indicates that $p\geq\mu_{1}$. Consequently, we have $p\geq\max\{p,q\}+r^{\prime}-m$, so we can discard the second inequality under the summation of (5.4). Additionally, $\max\{p,q\}+r^{\prime}-n=q+r^{\prime}-n\leq q$ because $r^{\prime}\leq\min\{m,n\}$. Thus, we can discard the last inequality under the summation of (5.4). Therefore, the claim above holds.
@@
The containment $I_{n,m}\subseteq I_{n,m,r}$ indicates an ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-equivariant surjection
@@
Then Lemma 2.2 and Corollary 3.9 indicates that
@@
However, Inequality (5.8) indicates that
@@ 6. Conclusion
We provide an open problem regarding log-concavity. Recall that a sequence of positive real numbers $(a_{1},a_{2},\cdots,a_{N})$ is log-concave if for all $1<i<N$ we have $a_{i}^{2}\geq a_{i-1}\cdot a_{i+1}$. Chen conjectured [1] that the sequence $(a_{n,1},a_{n,2},\cdots,a_{n,n})$ is log-concave where $a_{n,k}$ is the number of permutations in ${\mathfrak{S}}_{n}$ with longest subsequence of length $k$. Using the terminology of orbit harmonics, Chen’s conjecture means that the Hilbert series of the orbit harmonics ring $R({\mathfrak{S}}_{n})$ of the permutation matrix locus ${\mathfrak{S}}_{n}$ is log-concave.
We can generalize the definition of log-concavity by incorporating group action. Let $G$ be a group and $(V_{1},V_{2},\cdots,V_{N})$ is a sequence of $G$-modules. We say that $(V_{1},V_{2},\cdots,V_{N})$ is $G$-equivariant log-concave if for all $1<i<N$ we have a $G$-module surjection $V_{i}\otimes V_{i}\twoheadrightarrow V_{i-1}\otimes V_{i+1}$ where the $G$-action is diagonal, i.e., $g\cdot(v\otimes w)=(g\cdot v)\otimes(g\cdot w)$. A graded $G$-module $V=\bigoplus_{d=0}^{N}V_{d}$ is $G$-equivariant log-concave if the sequence $(V_{1},V_{2},\cdots,V_{N})$ is $G$-equivariant log-concave. Rhoades conjectured [13] that $R({\mathfrak{S}}_{n})$ is ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$-equivariant log-concave, which generalizes Chen’s conjecture.
@@
Another interesting future direction is to study the orbit harmonics ring $R({\mathcal{Z}}_{n,m,r})$. Since ${\mathcal{Z}}_{n,m,r}\subseteq{\mathcal{UZ}}_{n,m,r}$, we need to add some elements to the generating set in Definition 3.1 to obtain the defining ideal ${\mathrm{gr}}{\mathbf{I}}({\mathcal{Z}}_{n,m,r})$, such as $\sum_{i=1}^{n}\sum_{j=1}^{m}x_{i,j}$ and ${\mathfrak{m}}({\mathcal{R}})$ for ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r+1}$. Some relevant results will be presented in a forthcoming paper.

### v3

@@ 4.4. Standard monomial basis and Hilbert series
In order to prove Theorem 4.12, we need the following lemma on ideal containment.
Lemma 4.14.
$I_{n,m,r}\subseteq{\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$.
Proof.
For any rook placement ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$, each column (resp. row) of $n]\times[m]$ has at most one rook. Consequently, for all $1\leq i\leq n$ and $1\leq j,j^{\prime}\leq m$ we have $x_{i,j}\cdot x_{i,j^{\prime}}\in{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$ and thus $x_{i,j}\cdot x_{i,j^{\prime}}\in{\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$ (resp. for all $1\leq i,i^{\prime}\leq n$ and $1\leq j\leq m$ we have $x_{i,j}\cdot x_{i^{\prime},j}\in{\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$). Additionally, since $\mathcal{R}$} has at least $r$ rooks, any union of nr+1n-r+1 distinct columns of $n]\times[m]$ must meet $\mathcal{R}$}. Thus, for all $1\leq i_{1}<\dots<i_{n-r+1}\leq n$ we have
@@
Similarly, for all $1\leq j_{1}<\dots<j_{m-r+1}\leq m$ we have
@@
The proof is thus complete as all generators of $I_{n,m,r}$ are in ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{UZ}}_{n,m,r})$. ∎
With the lemma, we now give the proof of Theorem 4.12
Proof of Theorem 4.12.
Lemma 4.14yields a surjective $\mathbb{C}$}-linear map
@@
implying that $\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r})\geq\dim_{\mathbb{C}}(R({\mathcal{UZ}}_{n,m,r}))$. However, Lemma 4.10 implies that $\dim_{\mathbb{C}}({\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r})\leq|{\mathcal{UZ}}_{n,m,r}|=\dim_{\mathbb{C}}(R({\mathcal{UZ}}_{n,m,r}))$. Therefore, we conclude that
@@
where $c_{{\mathcal{R}}^{\prime}}\in{\mathbb{C}}$. However, Remark 4.11 implies that the right-hand side of Equation 4.2 can be replaced with a linear combination of some extended shadow monomials
@@
Theorem 4.12 immediately gives the Hilbert series of $R({\mathcal{UZ}}_{n,m,r})={\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$. Given a permutation $w\in{\mathfrak{S}}_{n+m-r}$, a increasing subsequence of length $k$ is a sequence $1\leq i_{1}<\dots<i_{k}\leq n+m-r$ of integers such that $w(i_{1})<\dots<w(i_{k})$. We write
@@
Let $k$ be the number of shadow lines of $w$ that we use to construct ${\mathcal{S}}(w)$. Write $w\mapsto(P,Q)$ according to the Schensted correspondence, where $P,Q$ have the same shape $\lambda$. The shadow line construction in Subsection 4.1 implies that $k=\lambda_{1}$. In addition, Schensted correspondence possesses the standard property [15, Thm. 1] $\lambda_{1}=\mathrm{lis}(w)$, so we have $k=\mathrm{lis}(w)$ and thus Equation (4.4) holds. ∎
@@ 5. Module structure
For $d,a,b,p,q\in{\mathbb{Z}}_{\geq 0}$, consider
$F:=\sum_{\mu\vdash d}\{s_{\mu}\cdot h_{a}\}_{\lambda_{1}=p}\otimes\{s_{\mu}\cdot h_{b}\}_{\lambda_{1}=q}\in\Lambda\otimes_{{\mathbb{C}}(q)}\Lambda$
@@
Let $d,a,b,p,q\in{\mathbb{Z}}_{\geq 0}$ be nonnegative integers. We have
@@
If one of them is not satisfied, Lemma 5.1 implies that $\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle F=0=\langle s_{\lambda^{(1)}}\otimes s_{\lambda^{(2)}}\rangle G$. We henceforth suppose that all these three conditions hold. Then Lemma 5.1 implies that
@@
Pieri’s rule and Lemma 5.2 imply that the left-hand side equals
@@
  • $p\geq n-d$ and $q\geq m-d$ together imply that $n+m-d-\max\{p,q\}\leq\min\{m,n\}$.
@@
Therefore, $0\leq\min\{m,n\}-r^{\prime}=n+m-\max\{m,n\}-r^{\prime}\leq n+m-\max\{p,q\}-r^{\prime}\leq n+m-(\max\{m,n\}-r^{\prime})-r^{\prime}=\min\{m,n\}$, so we have $0\leq n+m-\max\{p,q\}-r^{\prime}\leq\min\{m,n\}$ and hence the first equality under the summation in (5.4) can be discard. Now we suppose that $p\leq q$ without loss of generality. Then (5.6) implies that $\max\{p,q\}+r^{\prime}-m=q+r^{\prime}-m\leq\mu_{1}$, and (5.5) implies that $p\geq\mu_{1}$. Consequently, we have $p\geq\max\{p,q\}+r^{\prime}-m$, so we can discard the second inequality under the summation of (5.4). Additionally, $\max\{p,q\}+r^{\prime}-n=q+r^{\prime}-n\leq q$ because $r^{\prime}\leq\min\{m,n\}$. Thus, we can discard the last inequality under the summation of (5.4). Therefore, the claim above holds.
@@
The containment $I_{n,m}\subseteq I_{n,m,r}$ implies an ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$-equivariant surjection
@@
Then Lemma 2.2 and Corollary 3.9 implies that
@@
However, Inequality (5.8) implies that
@@ 6. Conclusion
We provide an open problem regarding log-concavity. Recall that a sequence of positive real numbers $(a_{1},a_{2},\dots,a_{N})$ is log-concave if for all $1<i<N$ we have $a_{i}^{2}\geq a_{i-1}\cdot a_{i+1}$. Chen conjectured [1] that the sequence $(a_{n,1},a_{n,2},\dots,a_{n,n})$ is log-concave where $a_{n,k}$ is the number of permutations in ${\mathfrak{S}}_{n}$ with longest subsequence of length $k$. Using the terminology of orbit harmonics, Chen’s conjecture means that the Hilbert series of the orbit harmonics ring $R({\mathfrak{S}}_{n})$ of the permutation matrix locus ${\mathfrak{S}}_{n}$ is log-concave.
We can generalize the definition of log-concavity by incorporating group action. Let $G$ be a group and $(V_{1},V_{2},\dots,V_{N})$ is a sequence of $G$-modules. We say that $(V_{1},V_{2},\dots,V_{N})$ is $G$-equivariant log-concave if for all $1<i<N$ we have a $G$-module surjection $V_{i}\otimes V_{i}\twoheadrightarrow V_{i-1}\otimes V_{i+1}$ where the $G$-action is diagonal, i.e., $g\cdot(v\otimes w)=(g\cdot v)\otimes(g\cdot w)$. A graded $G$-module $V=\bigoplus_{d=0}^{N}V_{d}$ is $G$-equivariant log-concave if the sequence $(V_{1},V_{2},\dots,V_{N})$ is $G$-equivariant log-concave. Rhoades conjectured [13] that $R({\mathfrak{S}}_{n})$ is ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$-equivariant log-concave, which generalizes Chen’s conjecture.
@@
Another interesting future direction is to study the orbit harmonics ring $R({\mathcal{Z}}_{n,m,r})$. Since ${\mathcal{Z}}_{n,m,r}\subseteq{\mathcal{UZ}}_{n,m,r}$, we need to add some elements to the generating set in Definition 3.1 to obtain the defining ideal ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}}_{n,m,r})$, such as $\sum_{i=1}^{n}\sum_{j=1}^{m}x_{i,j}$ and ${\mathfrak{m}}({\mathcal{R}})$ for ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r+1}$. Some relevant results are presented in Zhu [17] focusing on the graded module structure.